Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y = mx+b.

(-2,8; 9x = 5y + 2 The equation of the line is y= _____
Simplfy your answer. Use intergers for any numbers in the expression.

Answer by Maths68(1474)

Equation of the line slope-intercept form
y=mx+b
Given
Equation of the Given line
9x=5y+2
Rearrange above equation according to slope-intercept form
5y=9x-2
5y/5=(9x-2)/5
y=9/5(x)-2/5
Compare with the equation of line slope-intercept form
Slope of the given = m = 9/5
Since lines are parallel their slope will be same
Slope of the required line = m = 9/5
Point (-2, 8)
We have a point and slope we can get easily the require line equation.
Use Equation of the line point-slope form
m = y2-y1/x2-x1
Put the values in above equation, we have
9/5=(y-8)/(x-(-2))
9/5=(y-8)/(x+2)
9(x+2)=(y-8)5
9x+18=5y-40
9x+18+40=5y
9x+58=5y
(9x+58)/5=5y/5
9/5(x)+58/5=y
y=9/5(x)+58/5
Above equation is the equation of the required line in the form of y=mx+b